What is Zα/2 for a 90% confidence interval of the population mean? Group of answer choices 0.49 1.96 0.48 1.645  
Answer #1

Solution :

At 90% confidence level the z is ,

= 1 - 90% = 1 - 0.90 = 0.10

/ 2 = 0.10 / 2 = 0.05

Z/2 = Z 0.05 = 1.645